# RBI-ThPhys-2024-15

### v1

@@ Abstract
Interference effects are well founded from the quantum mechanical viewpoint and in principle cannot be ignored in realistic studies of New Physics scenarios. In this work, we investigate the size of interference effects between resonant and non-resonant contributions to di-Higgs production in the singlet extension of the Standard Model where the additional heavy scalar provides a resonant channel of on-shell and off-shell di-Higgs production. We find these interference contributions to have a non-negligible effect on the cross-sections and differential distributions. In order to allow for an computationally efficient treatment of these effects via reweighting, we introduce a new tool utilizing a matrix-element reweighting method: HHReweighter. In addition to the broadly used invariant di-Higgs mass $m_{hh}$, we analyze the sensitivity of other kinematic variables to the interference term and propose to also use differential distributions in the Higgs transverse momentum $p_{T}^{h}$. Furthermore, we provide updates on the latest experimental and theoretical limits on the parameter space of the real singlet extension of the Standard Model Higgs sector.
@@ I Motivation
Higgs pair production is a process of high interest at the LHC because of its sensitivity to the triple-Higgs coupling and therefore to the underlying mechanism of electroweak symmetry breaking and possible modifications of the scalar potential with respect to its prediction in the Standard Model (SM). Thus, a meaningful interpretation of the measured total and differential cross sections of this process requires an accurate theoretical description and simulation. We focus on di-Higgs production stemming from an additional resonance, as e.g. present in models with singlet extensions (see e.g. Robens:2015gla ; Robens:2016xkb ; Robens:2022cun and references therein). Di-Higgs production has been discussed in great detail in DiMicco:2019ngk (see also the short summary in Brigljevic:2024vuv ), where in particular finite widths and interference effects from Refs. Dawson:2015haa ; Dawson:2016ugw ; Lewis:2017dme ; Carena:2018vpt are presented. More recent work on interference effects in di-Higgs production in extended scalar sectors can be found e.g. in Refs. Basler:2019nas ; Heinemeyer:2024hxa . The assumption of using factorized cross sections has its own merits. One example is the fact that often it is much easier to calculate higher-order corrections to a single s-channel production and decay, without taking the full corrections into account. In some cases, these can be $\mathcal{O}\left(100\%\right)$, see e.g. Refs. Ellis:1996mzs ; Campbell:2017hsr , and might easily be more important than a leading-order calculation where the interference effects are correctly described. However, it is not clear a priori that this is the case in all scenarios and realizations of physical models. In particular, for some channels strong relations exist between couplings of new physics particles, e.g. via unitarity relations Gunion:1990kf or other theory constraints. In such cases, it is imperative to include the full matrix elements.
@@
Resonant di-scalar searches are currently investigated by both ATLAS and CMS ATLAS:2024ish ; CMS:2024phk , and correspond to one of the prime channels of testing the scalar sector realized in nature via triple scalar couplings. Therefore, it is of high importance that the correct theoretical description of such processes and use up-to-date tools are used in the corresponding simulations. For example, the correct interpretation of results relies on the correct understanding of the simulated process, that might include kinematic effects in differential distributions resulting from irreducible interference terms. Similarly, as the finite width of unstable particles is a model prediction it cannot be set to a value chosen independently of the involved masses and couplings.
@@
This paper is organized as follows. In Sect. II, we introduce the real singlet model used for our studies, and briefly discuss general features of width and interference effects in Sect. III. In Sect. IV and Sect. V we describe the tools used to simulate di-Higgs production and present our method for determining regions of the parameter space with sizable interference effects. We introduce our tool used to model interference effects using matrix-element reweighting in Sect. VI. This is followed by an investigation of interference effects for several benchmark points with distinct features in Sect. VII. We summarize our investigations in Sect. VIII.
@@ II.1 Summary of the model
$=$ $-m^{2}\Phi^{\dagger}\Phi-\mu^{2}S^{2}+\lambda_{1}(\Phi^{\dagger}\Phi)^{2}+\lambda_{2}S^{4}+\lambda_{3}\Phi^{\dagger}\Phi S^{2}.$ (8)
where we already imposed a $\mathbb{Z}_{2}$ symmetry that is softly broken by the vev of the singlet field $S$ [^1] In fact the general phenomenology and interference effects are independent of this choice and only depend on coupling rescaling and total width for a chosen parameter point. Applying symmetries has some impact on the available model parameter space, mainly through theoretical constraints..
After minimization, the model contains in total 5 free parameters in the electroweak sector. A physical choice is Robens:2015gla ; Robens:2016xkb
@@
where the masses of the two physical states are ordered as $m_{h}\,\leq\,m_{H}$, and where $v\,,v_{s}$ denote the vevs of the doublet $\Phi$ and singlet $S$, respectively. The mixing angle $\sin\alpha$ corresponds to the angle that rotates the gauge into the mass eigenstates,
@@
$m_{H},\,\sin\alpha,\,\tan\beta\,\equiv\,\frac{v}{v_{s}}$ (11)
@@
The trilinear Higgs couplings $\lambda_{ijk}$ (of mass dimension one) play a crucial role in the di-Higgs production processes. They are defined as $V\,\supset\,\sum_{ijk}\,\lambda_{ijk}\,h_{i}\,h_{j}\,h_{k}$, with $\left\{i,j,k\right\}\,\in\,\{1,2\}$ and $h_{1}\,\equiv\,h,\,h_{2}\,\equiv\,H$, where $\lambda_{hhh}$ is expressed in relation to the SM-value as $\kappa_{\lambda_{hhh}}\,\equiv\,\frac{\lambda_{hhh}}{\lambda_{\mathrm{SM}}}$, with $\lambda_{\mathrm{SM}}\,=\,\frac{m_{125}^{2}}{2\,v}$ being the triple scalar coupling in the SM like case.
@@
$\lambda_{hhh}$ $=$ $\lambda_{\mathrm{SM}}\left(\cos^{3}{\alpha}-\tan\beta\sin^{3}{\alpha}\right)$
$\lambda_{Hhh}$ $=$ $\lambda_{\mathrm{SM}}\frac{2m_{h}^{2}+m_{H}^{2}}{m_{h}^{2}}\,\frac{\sin\left(2\alpha\right)}{2}\left(\cos{\alpha}+\tan\beta\sin{\alpha}\right),$ (12)
@@
$\Gamma_{h}$ $=$ $\left(1-\sin^{2}\alpha\right)\Gamma_{\mathrm{SM}}$
@@
$\lambda_{hhh}$$\approx$ $\lambda_{\mathrm{SM}}\left(1-\frac{3}{2}\sin^{2}\alpha\right)$
@@ II.2 Constraints on the model
We include the most up-to-date experimental constraints. We do this by making use of the publicly available tool HiggsTools Bahl:2022igd . The allowed parameter space in the $\left(m_{H},\,\sin\alpha\right)$ and $\left(m_{H},\,\tan\beta\right)$ planes is shown in Fig. 1. Direct search results stem from $H\rightarrow\,Z\,Z$ CMS:2018amk ; ATLAS:2020tlo , $H\,\rightarrow\,V\,V$ ATLAS:2020fry , $H\,\rightarrow\,h\,h$ CMS:2021yci ; ATLAS:2021ifb ; ATLAS:2022xzm as well as combined searches ATLAS:2018sbw .
The parameter scan used here follows the prescription in Pruna:2013bma ; Robens:2015gla ; Robens:2016xkb ; Ilnicka:2018def . We first applied all possible theoretical and experimental constraints apart from direct search limits and signal strength measurements. This is due to the fact that most of these can be implemented directly in the scan in a straightforward way. The scan then also tests agreement with direct searches and signal strength using the standalone HiggsBounds and HiggsSignals version with results up to 2022. The allowed points are then checked against the most up to date constraints using the most recent version of HiggsTools222 Exclusion criteria within HiggsTools that are used here follow the prescription in Robens:2023pax , i.e. we use the observed and not the expected sensitivity. .
We show the most constraining channel, using the current HiggsTools version, from this last step in Fig. 1. This corresponds to an update of the results presented in previous work, as e.g. Robens:2015gla ; Robens:2016xkb ; Ilnicka:2018def ; Robens:2022cun .
Concerning the cross section values used for the determination of the rates, HiggsTools applies a factorized approach, where in particular the production cross sections are obtained by rescaling the respective values given in the Yellow Report 4 of the Higgs working group LHCHiggsCrossSectionWorkingGroup:2016ypw , rescaled by $\sin^{2}\,\alpha$ and $\cos^{2}\alpha$ for $H$ and $h$, respectively. Note that care must be taken to distinguish the validity regimes of the respective calculations: while the N3LO calculation relies on the effective theory and is therefore only viable for $m_{H}\,\leq\,m_{\text{top}}$, for higher masses the NNLO+NNLL value must be taken which contains full top quark mass dependence at leading order.
Figure 1: Allowed parameter space in the real singlet extension in the $\left(m_{H},\,\sin\alpha\right)$ (left) and $\left(m_{H},\,\tan\beta\right)$ (right) planes. Maximal values for the third free parameter are shown in color coding. In addition, on the left side we also display exclusion bounds from the $W$ mass measurement (purple), as well as several resonance searches at the LHC (color-shaded areas). See text for details.
@@ III.1 Narrow width approximation
In general, from a quantum field theoretical viewpoint, only stable particles can be used as the incoming and outgoing states of the S-matrix (see e.g. Ref. Itzykson:1980rh ), implying that all unstable particles can be treated only as intermediate states (see also the discussion in e.g. Denner:2019vbn ). In practise, however, it can be advantageous to make use of the narrow width approximation (NWA) pilkuhn1967interactions ; Dicus:1984fu to factorize a more complicated process into the on-shell production of an intermediate particle and its subsequent decay. This way, one can conveniently include higher-order corrections to the production and decay processes. We briefly review the assumptions underlying this approximation.
Let us consider a process $a\,b\,\rightarrow\,S\,\rightarrow\,c\,d$, where we now assume $S$ can be produced and decay resonantly and $a,\,b,\,c,\,d$ are stable initial and final state particles, respectively. In the case the $s$-channel is the only contributing diagram, the matrix element can be written via the resummation of the infinite series of self-energy insertions $\Sigma(p^{2})$ between the tree-level propagators $D(p^{2})=i/(p^{2}-m_{S}^{2})$ (see Denner:2019vbn ) as
@@
In Eq. (16), we used the optical theorem to relate the quantum field theoretical renormalized self-energy to the physical total width of $S$, $\Gamma_{S}$.
$\text{Im}\hat{\Sigma}\,\left(p^{2}\,=\,m_{S}^{2}\right)\,=\,\Gamma_{S}\,m_{S}$ (17)
at the pole $p^{2}=m_{S}^{2}$, giving rise to the well-known Breit-Wigner propagator in the final step of Eq. (16); for a detailed discussion see e.g. Ref. Fuchs:2016swt . The total width $\Gamma$ corresponds to the sum over all possible partial decay widths [^3] We here for simplicity assume that $m_{S}$ is real. To ensure gauge invariance, it is helpful to work in the complex mass scheme, see e.g. Denner:1999gp ; Denner:2005fg ; Denner:2006ic ; Denner:2014zga ; Denner:2019vbn for details. .
@@
One can also use the Breit-Wigner propagator on the left-hand side of Eq. (18) for finite widths. Further assumptions in the NWA are that the function $\mathcal{F}$ only obtains major contributions around $p^{2}\,\sim\,m^{2}$ and only varies mildly in the region $p^{2}\,\in\,\left[\left(m-\Gamma\right)^{2},\left(m+\Gamma\right)^{2}\right]$ [^4] This assumption, and the validity of the NWA, can be violated e.g. near kinematic thresholds.. In this case, and making use of phase space factorization, one obtains the $s$-channel mediated cross section
@@ III.2 Interference effects
In case the process is dominated by an on-shell resonance $S$, very often the process is well-described by the first contribution in the above equation. However, it is not a priori clear that this holds for all kinematic variables in the process, including differential distributions or specific regions of phase space. In such cases, all above contributions need to be taken into account. In particular, for processes with electroweak vector bosons in the initial and final state that include scalar $s$-channel contributions, unitarity links these Gunion:1990kf and requires that the full squared matrix element is included in a correct description.

### v2

@@ Abstract
Interference effects are well founded from the quantum mechanical viewpoint and in principle cannot be ignored in realistic studies of New Physics scenarios. In this work, we investigate the size of interference effects between resonant and non-resonant contributions to di-Higgs production in the singlet extension of the Standard Model, where the additional heavy scalar provides a resonant channel. We find these interference contributions to have a non-negligible effect on the cross-sections and differential distributions. In order to allow for a computationally efficient treatment of these effects via reweighting, we introduce a new tool utilising a matrix-element reweighting method: HHReweighter. In addition to the broadly used di-Higgs invariant mass $m_{hh}$, we analyse the sensitivity to the interference terms for other kinematic variables, such as the Higgs boson transverse momentum, and find that these also can be sensitive to interference effects. Furthermore, we provide updates on the latest experimental and theoretical limits on the parameter space of the real singlet extension of the Standard Model Higgs sector.
@@ I Motivation
Higgs pair production is a process of high interest at the LHC because of its sensitivity to the triple-Higgs coupling and therefore to the underlying mechanism of electroweak symmetry breaking and possible modifications of the scalar potential with respect to its prediction in the Standard Model (SM). Thus, a meaningful interpretation of the measured total and differential cross sections of this process requires an accurate theoretical description and simulation. We focus on di-Higgs production stemming from an additional resonance, as e.g. present in models with singlet extensions (see e.g. Refs. Robens:2015gla ; Robens:2016xkb ; Robens:2022cun and references therein). Di-Higgs production has been discussed in great detail in Ref. DiMicco:2019ngk (see also the short summary in Ref. Brigljevic:2024vuv ), where in particular finite widths and interference effects from Refs. Dawson:2015haa ; Dawson:2016ugw ; Lewis:2017dme ; Carena:2018vpt are presented. More recent work on interference effects in di-Higgs production in extended scalar sectors can be found e.g. in Refs. Basler:2019nas ; Heinemeyer:2024hxa .
The assumption of using factorised cross sections has its own merits. One example is the fact that often it is much easier to calculate higher-order corrections to a single $s$-channel production and decay, without taking the full corrections into account. In some cases, these can be $\mathcal{O}\left(100\%\right)$, see e.g. Refs. Ellis:1996mzs ; Campbell:2017hsr , and might easily be more important than a leading order (LO) calculation where the interference effects are correctly described. However, it is not clear a priori that this is the case in all scenarios and realisations of physical models. In particular, for some channels strong relations exist between couplings of new physics particles, e.g. via unitarity relations Gunion:1990kf or other theory constraints. In such cases, it is imperative to include the full matrix elements.
@@
Resonant di-scalar searches are currently investigated by both ATLAS and CMS ATLAS:2024ish ; CMS:2024phk , and correspond to one of the prime channels for testing the scalar sector realised in nature via triple scalar couplings. Therefore, it is of high importance that the correct theoretical description of such processes is consistently applied and up-to-date tools are used in the corresponding simulations. For example, the correct interpretation of results relies on the correct understanding of the simulated process, that might include kinematic effects in differential distributions resulting from irreducible interference terms. First experimental studies showing the possible importance of interference effects in di-Higgs production were presented by the CMS Collaboration in Ref. CMS:2024phk . Similarly, as the finite width of unstable particles is a model prediction, it cannot be set to a value chosen independently of the involved masses and couplings. Note that modifications to the di-Higgs kinematic distributions can also appear in effective field theory interpretations, see e.g. Refs. Carvalho:2015ttv ; Buchalla:2018yce ; Capozi:2019xsi ; Alasfar:2023xpc .
@@
This paper is organised as follows. In Section II, we introduce the real singlet model used for our studies, and briefly discuss general features of width and interference effects in Section III. In Section IV and Section V we describe the tools used to simulate di-Higgs production and present our method for determining regions of the parameter space with sizeable interference effects. We introduce our tool used to model interference effects using matrix-element reweighting in Section VI. This is followed by an investigation of interference effects for several benchmark points with distinct features in Section VII. We summarise our investigations in Section VIII.
@@ II.1 Summary of the model
$=$ $-m^{2}\Phi^{\dagger}\Phi-\mu^{2}S^{2}+\lambda_{1}(\Phi^{\dagger}\Phi)^{2}+\lambda_{2}S^{4}+\lambda_{3}\Phi^{\dagger}\Phi S^{2},$ (8)
where we already imposed a $\mathbb{Z}_{2}$ symmetry that is softly broken by the vacuum expectation value (VEV) of the singlet field $S$ [^1] In fact the general phenomenology and interference effects are independent of this choice and only depend on the coupling rescaling and the total width for a chosen parameter point. Applying symmetries has some impact on the available model parameter space, mainly through theoretical constraints..
After minimisation, the model contains in total 5 free parameters in the electroweak sector. A physical choice is Robens:2015gla ; Robens:2016xkb
@@
where the masses of the two physical states are ordered as $m_{h}\,\leq\,m_{H}$, and where $v$ and $v_{s}$ denote the VEVs of the doublet $\Phi$ and singlet $S$, respectively. The mixing angle $\sin\alpha$ corresponds to the angle that rotates the gauge into the mass eigenstates,
@@
$m_{H},\,\sin\alpha,\,\tan\beta\,\equiv\,\frac{v}{v_{s}},$ (11)
@@
The trilinear Higgs couplings $\lambda_{ijk}$ (of mass dimension one) play a crucial role in the di-Higgs production processes. They are defined as $V\,\supset\,\sum_{ijk}\,\lambda_{ijk}\,h_{i}\,h_{j}\,h_{k}$, with $\left\{i,j,k\right\}\,\in\,\{1,2\}$ and $h_{1}\,\equiv\,h,\,h_{2}\,\equiv\,H$, where $\lambda_{hhh}$ is expressed in relation to the SM-value as $\kappa_{\lambda_{hhh}}\,\equiv\,\frac{\lambda_{hhh}}{\lambda_{\mathrm{SM}}}$, with $\lambda_{\mathrm{SM}}\,=\,\frac{m_{125}^{2}}{2\,v}$ being the triple scalar coupling in the SM-like case.
@@
$\lambda_{hhh}$ $=$ $\lambda_{\mathrm{SM}}\left(\cos^{3}{\alpha}-\tan\beta\sin^{3}{\alpha}\right),$
$\lambda_{Hhh}$ $=$ $\lambda_{\mathrm{SM}}\frac{2m_{h}^{2}+m_{H}^{2}}{m_{h}^{2}}\,\frac{\sin\left(2\alpha\right)}{2}\left(\cos{\alpha}+\tan\beta\sin{\alpha}\right).$ (12)
@@
$\Gamma_{h}$ $=$ $\left(1-\sin^{2}\alpha\right)\Gamma_{\mathrm{SM}},$
@@
$\lambda_{hhh}$$\approx$ $\lambda_{\mathrm{SM}}\left(1-\frac{3}{2}\sin^{2}\alpha\right),$
@@ II.2 Constraints on the model
We include the most up-to-date experimental constraints. We do this by making use of the publicly available tool HiggsTools Bahl:2022igd , that builds on HiggsBounds Bechtle:2008jh ; Bechtle:2011sb ; Bechtle:2013wla ; Bechtle:2020pkv and HiggsSignals Bechtle:2013xfa ; Bechtle:2014ewa ; Bechtle:2020uwn . The allowed parameter space in the $\left(m_{H},\,\sin\alpha\right)$ and $\left(m_{H},\,\tan\beta\right)$ planes is shown in Figure 1. Direct search results stem from $H\rightarrow\,Z\,Z$ CMS:2018amk ; ATLAS:2020tlo , $H\,\rightarrow\,V\,V$ ATLAS:2018sbw ; ATLAS:2020fry , and $H\,\rightarrow\,h\,h$ CMS:2021yci ; ATLAS:2021ifb ; ATLAS:2022xzm .
The parameter scan used here follows the prescription in Refs. Pruna:2013bma ; Robens:2015gla ; Robens:2016xkb ; Ilnicka:2018def . We first applied all possible theoretical and experimental constraints apart from direct search limits and signal strength measurements. This is due to the fact that most of these can be implemented directly in the scan in a straightforward way. The scan then also tests agreement with direct searches and signal strength measurements using the standalone HiggsBounds and HiggsSignals version with results up to 2022. The allowed points are then checked against the most up to date constraints using the most recent version of HiggsTools 222 Exclusion criteria within HiggsTools that are used here follow the prescription in Ref. Robens:2023pax , i.e. we use the observed and not the expected sensitivity. . The $W$-boson mass was calculated according to Refs. Lopez-Val:2014jva ; Papaefstathiou:2022oyi .
We show the most constraining channel, using the current HiggsTools version, from this last step in Figure 1. This corresponds to an update of the results presented in previous work, as e.g. Robens:2015gla ; Robens:2016xkb ; Ilnicka:2018def ; Robens:2022cun .
Concerning the cross section values used for the determination of the rates, HiggsTools applies a factorised approach, where the production cross sections are obtained by rescaling the respective values given in the Yellow Report 4 of the Higgs working group LHCHiggsCrossSectionWorkingGroup:2016ypw , rescaled by $\sin^{2}\,\alpha$ and $\cos^{2}\alpha$ for $H$ and $h$, respectively. Note that care must be taken to distinguish the validity regimes of the respective calculations: for $m_{H}\,>\,m_{\text{top}}$ the next-to-next-to-LO (NNLO) with next-to-next-to-leading logarithmic resummation value must be taken, as it contains full top quark mass dependence up to next-to-LO order (NLO), while the next-to-NNLO calculation relies on the effective theory and is therefore viable for $m_{H}\,\leq\,m_{\text{top}}$.
Figure 1: Allowed parameter space in the real singlet extension in the $\left(m_{H},\,\sin\alpha\right)$ (left) and $\left(m_{H},\,\tan\beta\right)$ (right) planes. Maximal values for the third free parameter are shown in colour coding. In addition, on the left side we also display exclusion bounds from the $W$ mass measurement (purple), as well as several resonance searches at the LHC (colour-shaded areas). See text for details.
@@ III.1 Narrow width approximation
In general, from a quantum field theoretical viewpoint, only stable particles can be used as the incoming and outgoing states of the S-matrix (see e.g. Ref. Itzykson:1980rh ), implying that all unstable particles can be treated only as intermediate states (see also the discussion in e.g. Ref. Denner:2019vbn ). In practise, however, it can be advantageous to make use of the narrow width approximation (NWA) pilkuhn1967interactions ; Dicus:1984fu to factorise a more complicated process into the on-shell production of an intermediate particle and its subsequent decay. This way, one can conveniently include higher-order corrections to the production and decay processes. We briefly review the assumptions underlying this approximation.
Let us consider a process $a\,b\,\rightarrow\,S\,\rightarrow\,c\,d$, where we now assume $S$ can be produced and decay resonantly and $a,\,b,\,c,$ and $d$ are stable initial and final state particles, respectively. In the case the $s$-channel is the only contributing diagram, the matrix element can be written via the resummation of the infinite series of self-energy insertions $\Sigma(p^{2})$ between the tree-level propagators $D(p^{2})=i/(p^{2}-m_{S}^{2})$ (see Denner:2019vbn ) as
@@
In Eq. (16), we used the optical theorem to relate the quantum field theoretical renormalized self-energy, $\hat{\Sigma}$, to the physical total width of $S$, $\Gamma_{S}$, at the pole $p^{2}=m_{S}^{2}$:
$\text{Im}\hat{\Sigma}\,\left(p^{2}\,=\,m_{S}^{2}\right)\,=\,\Gamma_{S}\,m_{S},$ (17)
giving rise to the well-known Breit-Wigner propagator in the final step of Eq. (16). For a detailed discussion see e.g. Ref. Fuchs:2016swt . The total width $\Gamma$ corresponds to the sum over all possible partial decay widths [^3] We here for simplicity assume that $m_{S}$ is real. To ensure gauge invariance, it is helpful to work in the complex mass scheme, see e.g. Denner:1999gp ; Denner:2005fg ; Denner:2006ic ; Denner:2014zga ; Denner:2019vbn for details. .
@@
One can also use the Breit-Wigner propagator on the left-hand side of Eq. (18) for finite widths. Further assumptions in the NWA are that the function $\mathcal{F}$ only obtains major contributions around $p^{2}\,\sim\,m^{2}$ and only varies mildly in the region $p^{2}\,\in\,\left[\left(m-\Gamma\right)^{2},\left(m+\Gamma\right)^{2}\right]$ [^4] This assumption, and the validity of the NWA, can be violated e.g. near kinematic thresholds.. In this case, and making use of phase space factorisation, one obtains the $s$-channel mediated cross section
@@ III.2 Interference effects
In case the process is dominated by an on-shell resonance $S$, very often the process is well-described by the first contribution in the above equation. However, it is not a priori clear that this holds for all kinematic variables in the process, including differential distributions or specific regions of phase space. In such cases, all above contributions need to be taken into account. In particular, for processes with electroweak vector bosons in the initial and final state that include scalar $s$-channel contributions, unitarity links these Gunion:1990kf and requires that the full squared matrix element is included for a correct description.